### v1

@@ 4 Generalized NR-CDT
$\langle\mathcal{R}_{\phi,\gamma}[\mu],g\rangle=\langle\mu,\mathcal{R}_{\phi,\gamma}^{*}[g]\rangle\quad\forall\,g\in L_{\lambda_{\mathbb{R}}\times\gamma}^{\infty}(\mathbb{R}\times\Theta).$
@@
For all $\mu\in\mathcal{M}(\mathbb{X})$ and $g\in L^{\infty}(\mathbb{R}\times\Theta)$, applying Fubini’s theorem gives
@@
To introduce a generalized $h_{\mathrm{}}$NR-CDT, we have to ensure that each normalization step is well-defined. For this, we restrict ourselves to uniformly bounded defining functions $\phi$ meaning that, for every $L>0$, there exits $M>0$ satisfying
$\lvert\phi(\mathbf{x},{\bm{\theta}})\rvert\leq M\qquad\forall\,\lVert\mathbf{x}\rVert\leq L,~\forall\,{\bm{\theta}}\in\Theta.$
Furthermore, we only consider measures from
@@
By construction, $\mathcal{N}_{\phi}[\mu]$ is bounded on $\operatorname{supp}(\rho)\times\Theta$. In the style of § 3.4, let $h\colon L^{\infty}(\Theta)\to\mathbb{R}$ be bounded. Unless otherwise stated, we assume that $h(g\circ\psi)=h(g)$ for all $g\in L^{\infty}(\Theta)$ and any bijection $\psi\colon\Theta\to\Theta$, which is a stricter setting as in § 3.4 to allow for a general discussion. For $\mu\in\mathcal{P}_{\phi,\mathrm{c}}^{+}(\mathbb{X})$, the generalized $h_{\mathrm{}}$NR-CDT $\mathcal{N}_{h,\phi}[\mu]\colon\mathbb{R}\to\mathbb{R}$ may be defined via
@@ 4.2 Circular NR-CDT
Figuratively, the circular Radon transform integrates along circles with center ${\bm{\theta}}\in\Theta$. The resulting $h_{\mathrm{}}$NR-CDT is especially useful for classification tasks under isotropic affine transformations, i.e., for probability classes build by
@@
Because of (9), we have $\mathcal{N}_{\phi,h}[\mathbb{F}]=\bigl\{\mathcal{N}_{\phi,h}[\mu_{0}]\bigr\}$ and $\mathcal{N}_{\phi,h}[\mathbb{G}]=\bigl\{\mathcal{N}_{\phi,h}[\nu_{0}]\bigr\}$. Then $\mathcal{N}_{\phi,h}[\mu_{0}]\neq\mathcal{N}_{\phi,h}[\nu_{0}]$ implies their linear separability in $L^{\infty}_{\rho}(\mathbb{R})$. ∎
@@ 5 Numerical Experiments
Before reporting our numerical results, we wish to comment on some details regarding the numerical implementation. In the two-dimensional setting, each image is modelled as a piecewise constant function with support in $\tfrac{1}{\sqrt{2}},\tfrac{1}{\sqrt{2}}]^{2}\subset\mathbb{B}_{2}$, which allows us to compute its Radon transform in closed form. For numerical stability, we replace the line integrals in (1) by integrals over disjoint stripes of width $\frac{2}{R}$, where $R$ is the number of radial samples. Thereon, the CDTs are computed based on linear interpolation on an equispaced grid of interpolation points in $(0,1)$. In the multi-dimensional setting, we model the objects as empirical measures on $\mathbb{X}$. Then, for a discretized direction set $\Theta$, the restricted Radon transforms in (7) can be explicitly computed by evaluating (6), leading to piecewise constant functions on $[0,1]$. Then, the CDTs are exactly calculated by explicit inversion.
5.1 Image Classification
  class 1     class 2     class 3
$\mathrm{m}$NR-CDT $h_{\mathrm{a}}$NR-CDT $h_{\mathrm{b}}$NR-CDT
Refer to caption Refer to caption Refer to caption
$h_{\mathrm{c}}$NR-CDT $h_{\mathrm{d}}$NR-CDT $\mathrm{tv}$NR-CDT
Refer to caption Refer to caption Refer to caption
template 1 template 2 template 3
Refer to caption Refer to caption Refer to caption
Our two-dimensional numerical experiments make use of the following three datasets:
@@ 5.1.1 Nearest Template Classification
angles Eucl. R-CDT $\mathrm{m}$NR-CDT $h_{\mathrm{a}}$NR-CDT $h_{\mathrm{b}}$NR-CDT $h_{\mathrm{c}}$NR-CDT $h_{\mathrm{d}}$NR-CDT $\mathrm{tv}$NR-CDT
$\|\cdot\|_{\infty}$ $\|\cdot\|_{2}$ $\|\cdot\|_{\infty}$ $\|\cdot\|_{2}$ $\|\cdot\|_{\infty}$ $\|\cdot\|_{2}$ $\|\cdot\|_{\infty}$ $\|\cdot\|_{2}$ $\|\cdot\|_{\infty}$ $\|\cdot\|_{2}$ $\|\cdot\|_{\infty}$ $\|\cdot\|_{2}$ $\|\cdot\|_{\infty}$ $\|\cdot\|_{2}$ $\|\cdot\|_{\infty}$ $\|\cdot\|_{2}$
2 0.333 0.366 0.466 0.433 0.400 0.466 0.333 0.366 0.400 0.500 0.200 0.333 0.133 0.333 0.133 0.333
@@
angles Eucl. R-CDT $\mathrm{m}$NR-CDT $h_{\mathrm{d}}$NR-CDT $\mathrm{tv}$NR-CDT
2 0.144 0.122 0.089 0.122 0.122
@@
In order to observe the predicted behaviour numerically, the underlying Radon transform and CDT have to be discretized fine enough. Therefor, we choose different numbers of equispaced angles and $850$ equispaced radii for the Radon transform as well as $256$ equispaced interpolation points for the CDT.
We start with reporting our NT results for the academic dataset with $10$ samples per class, where we compare the Chebychev norm $\|\cdot\|_{\infty}$ and Euclidean norm $\|\cdot\|_{2}$ in $h_{\mathrm{}}$NR-CDT space for determining the nearest template. For illustration, the $h_{\mathrm{}}$NR-CDTs of the entire dataset are depicted in Figure 5, where the thick lines correspond to the templates and the thin lines to the transformed images. As predicted by our theory, the lines per class are nearly perfectly aligned. This also holds for the $\mathrm{tv}$NR-CDT, which in theory requires more regularity of the measures. An inspection of Figure 6, however, reveals that the corresponding NR-CDTs of the templates are rather smooth.
Our classification results are shown in Table 2 for varying numbers of equispaced angles for the underlying Radon transform. We observe that the $h_{\mathrm{}}$NR-CDT feature representations clearly outperform the R-CDT and Euclidean baseline and, remarkably, the classification is already perfect for a small number of chosen angles. As the different $h_{\mathrm{}}$NR-CDT representations perform nearly on par, we henceforth consider only $\mathrm{m}$NR-CDT, $h_{\mathrm{d}}$NR-CDT and $\mathrm{tv}$NR-CDT. Moreover, in most cases, $\|\cdot\|_{2}$ outperforms $\|\cdot\|_{\infty}$ so that from now on we restrict ourselves to the Euclidean norm $\|\cdot\|_{2}$.
To study the robustness of the NT classification we consider the polygon dataset with $10$ samples per class, which includes small non-affine perturbations. Hence, a perfect classification according to Theorems 1 and 2 cannot be expected. The accuracy results are reported in Table 3. While the Euclidean baseline and R-CDT are as bad as random guessing, $\mathrm{m}$NR-CDT and $h_{\mathrm{d}}$NR-CDT perform significantly better and, for sufficiently many Radon angles, $\mathrm{tv}$NR-CDT even reaches perfect results. This indicates its potential in real-world applications and may be attributed to the fact that $\mathrm{tv}$NR-CDT uses the whole information over the entire range of angles in $0,2\pi)$, while $\mathrm{m}$NR-CDT and $h_{\mathrm{d}}$NR-CDT only consider the largest and smallest angle information.

### v2

@@ 4 Generalized NR-CDT
$\langle\mathcal{R}_{\phi,\gamma}[\mu],g\rangle=\langle\mu,\mathcal{R}_{\phi,\gamma}^{*}[g]\rangle\quad\forall\,g\in\mathrm{L}_{\lambda_{\mathbb{R}}\times\gamma}^{\infty}(\mathbb{R}\times\Theta).$
@@
For all $\mu\in\mathcal{M}(\mathbb{X})$ and $g\in\mathrm{L}^{\infty}(\mathbb{R}\times\Theta)$, applying Fubini’s theorem gives
@@
To introduce a generalized $h_{\mathrm{}}$NR-CDT, we have to ensure that each normalization step is well-defined. For this, we restrict ourselves to uniformly bounded defining functions $\phi$ meaning that, for every $L>0$, there exits $M_{L}>0$ satisfying
$\lvert\phi(\mathbf{x},{\bm{\theta}})-\phi(\mathbf{y},{\bm{\theta}})\rvert\leq M_{L}\kern 5.0pt\forall\,d_{\mathbb{X}}(\mathbf{x},\mathbf{y})\leq L,~\forall\,{\bm{\theta}}\in\Theta,$
where $d_{\mathbb{X}}$ denotes a topology-compatible metric on $\mathbb{X}$. Furthermore, we only consider measures from
@@
By construction, $\mathcal{N}_{\phi}[\mu]$ is bounded on $\operatorname{supp}(\rho)\times\Theta$. In the style of § 3.4, let $h\colon\mathrm{L}^{\infty}(\Theta)\to\mathbb{R}$ be bounded. Unless otherwise stated, we assume that $h(g\circ\psi)=h(g)$ for all $g\in\mathrm{L}^{\infty}(\Theta)$ and any bijection $\psi\colon\Theta\to\Theta$, which is a stricter setting as in § 3.4 to allow for a general discussion. For $\mu\in\mathcal{P}_{\phi,\mathrm{c}}^{+}(\mathbb{X})$, the generalized $h_{\mathrm{}}$NR-CDT $\mathcal{N}_{h,\phi}[\mu]\colon\mathbb{R}\to\mathbb{R}$ may be defined via
@@ 4.2 Circular NR-CDT
Note that, due to the reversed triangle inequality, this function satisfies the above uniform boundedness assumption. Figuratively, the circular Radon transform integrates along circles with center ${\bm{\theta}}\in\Theta$. The resulting $h_{\mathrm{}}$NR-CDT is especially useful for classification tasks under isotropic affine transformations, i.e., for probability classes build by
@@
Because of (11), we have $\mathcal{N}_{\phi,h}[\mathbb{F}]=\bigl\{\mathcal{N}_{\phi,h}[\mu_{0}]\bigr\}$ and $\mathcal{N}_{\phi,h}[\mathbb{G}]=\bigl\{\mathcal{N}_{\phi,h}[\nu_{0}]\bigr\}$. Then $\mathcal{N}_{\phi,h}[\mu_{0}]\neq\mathcal{N}_{\phi,h}[\nu_{0}]$ implies their linear separability in $\mathrm{L}^{\infty}_{\rho}(\mathbb{R})$. ∎
@@ 5 Numerical Experiments
Before reporting our numerical results, we wish to comment on some details regarding the numerical implementation. In the two-dimensional setting, each grayscale image is modeled as a non-negative piecewise constant function $f\in\mathrm{L}^{1}(\mathbb{R}^{2})$ with support in $1/\sqrt{2},1/\sqrt{2}]^{2}\subset\mathbb{B}_{2}$. Moreover, we normalize $f$ such that $\int_{\mathbb{B}_{2}}f(\mathbf{x})\mathop{}\!\mathrm{d}\mathbf{x}=1$ to be interpretable as a probability density function. This allows us to compute its Radon transform in closed form. For numerical stability, we replace the line integrals in (1) by integrals over disjoint stripes of width $\frac{2}{R}$, where $R$ is the number of radial samples. Thereon, the CDTs are computed based on linear interpolation on an equispaced grid of interpolation points in $(0,1)$. In the multi-dimensional setting, we model the objects as empirical measures on $\mathbb{X}$. Then, for a discretized direction set $\Theta$, the restricted Radon transforms in (9) are themselves empirical measures that can be explicitly computed by evaluating (8). Consequently, the CDTs are piecewise constant functions and can be calculated exactly.
Algorithm 1 Generalized NR-CDT
1: input probability measure $\mu\in\mathcal{P}_{\phi,\mathrm{c}}^{+}(\mathbb{X})$, $\triangleright$ image / empirical measure
Radon directions ${\bm{\theta}}_{1},\dots,{\bm{\theta}}_{n}\in\Theta$,
number of sampling points $m\in\mathbb{N}$, aggregation map $h\colon\mathbb{R}^{m}\to\mathbb{R}$
2: $\mathbf{a}=\mathcal{N}_{\phi,h}[\mu]$
3: for $k=1,\dots,n$ do
4:   $\nu\leftarrow\mathcal{R}_{\phi,{\bm{\theta}}_{k}}[\mu]$ $\triangleright$ sampled density / empirical measure
5:   $\mathbf{q}\leftarrow\bigl[F^{[-1]}_{\nu}\bigl(\tfrac{1}{m+1}\bigr),\dots,F^{[-1]}_{\nu}\bigl(\tfrac{m}{m+1}\bigr)\bigr]$ $\triangleright$ linear interpolation / explicit inversion
6:   $\eta\leftarrow\tfrac{1}{m}\sum_{\ell=1}^{m}\mathbf{q}_{\ell}$
7:   $\sigma\leftarrow\sqrt{\tfrac{1}{m}\sum_{\ell=1}^{m}(\mathbf{q}_{\ell}-\eta)^{2}}$
8:   $\mathbf{N}_{\bullet,k}\leftarrow(\mathbf{q}\ominus\eta)\oslash\sigma$ $\triangleright$ pointwise subtraction & division
9: for $\ell=1,\dots,m$ do
10:   $\mathbf{a}_{\ell}=h(\mathbf{N}_{\ell,\bullet})$
[…]
For the uniform reference measure $\rho=u_{[0,1]}$, the computation of the CDT basically consists of calculating the cumulative sums of the $M$ Radon samples, which can be combined with the interpolation step. Hence, for $m$ CDT sampling points, we can calculate the CDT with $\mathcal{O}(M+m)$ operations. Computation of mean and standard deviation as well as the normalization step require $\mathcal{O}(m)$ operations. Finally, if $n$ is the number of Radon directions, all our considered aggregation functions have a complexity of $\mathcal{O}(n)$. Hence, the total computational complexity of the generalized NR-CDT is of order $\mathcal{O}(n(MN+m))$ for two-dimensional images and of order $\mathcal{O}(n(dN+m))$ for $d$-dimensional point clouds. Especially for images, where $N$ is usually large, the complexity can be lowered by exploiting the Fourier slice theorem and the non-equispaced fast Fourier transform.
5.1 Image Classification
@@
Our two-dimensional numerical experiments make use of the following four datasets:
@@
    agnus dei bishop crown lilies (1st) pinecone keys lilies (2nd) moon poppy orb shield
    Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
    Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
    Figure 5: Selection of three independent samples per class from the historic watermark dataset. Here, black color corresponds to positive mass as opposed to Fig. 24.
  4. d)
    Historic watermark dataset. Inspired by our motivating application in filigranology, we extend the historic watermark experiment from [37, § 5.5]. More precisely, we select eleven different classes from the watermark information system (WZIS) [^1] https://www.wasserzeichen-online.de, each consisting of 4–13 samples. The resulting dataset comprises 75 images with $256\times 256$ pixels, which are manually pre-processed by removing artificial vertical lines coming from the imaging technique. For illustration, three samples of each class are visualized in Figure 5.
@@ 5.1.1 Nearest Template Classification
  class 1     class 2     class 3
$\mathrm{m}$NR-CDT $h_{\mathrm{a}}$NR-CDT $h_{\mathrm{b}}$NR-CDT
Refer to caption Refer to caption Refer to caption
$h_{\mathrm{c}}$NR-CDT $h_{\mathrm{d}}$NR-CDT $\mathrm{tv}$NR-CDT
Refer to caption Refer to caption Refer to caption
In order to observe the predicted behaviour numerically, the underlying Radon transform and CDT have to be discretized fine enough. Therefore, we choose different numbers of equispaced angles and $850$ equispaced radii for the Radon transform as well as $256$ equispaced interpolation points for the CDT. In all our examples, the grid for the Radon radii is at least three times finer than the pixel size. Consequently, the considered images cannot be concentrated on a single Radon line, i.e., they are contained in $\mathcal{P}_{\mathrm{c}}^{*}(\mathbb{R}^{2})$.
We start with reporting our NT results for the academic dataset with $10$ samples per class, where we compare the Chebychev norm $\|\cdot\|_{\infty}$ and Euclidean norm $\|\cdot\|_{2}$ in $h_{\mathrm{}}$NR-CDT space for determining the nearest template. We compare our proposed feature extractors with the Euclidean baseline, a moment-based (MB) feature extractor [47], and linearized optimal transport (LOT) [9, 7].
template 1 template 2 template 3
Refer to caption Refer to caption Refer to caption
angles Eucl. MB LOT R-CDT $\mathrm{m}$NR-CDT $h_{\mathrm{a}}$NR-CDT $h_{\mathrm{b}}$NR-CDT $h_{\mathrm{c}}$NR-CDT $h_{\mathrm{d}}$NR-CDT $\mathrm{tv}$NR-CDT
$\nicefrac{{\|\cdot\|_{\infty}}}{{\|\cdot\|_{2}}}$ $\nicefrac{{\|\cdot\|_{\infty}}}{{\|\cdot\|_{2}}}$ $\|\cdot\|_{\infty}$ $\|\cdot\|_{2}$ $\|\cdot\|_{\infty}$ $\|\cdot\|_{2}$ $\|\cdot\|_{\infty}$ $\|\cdot\|_{2}$ $\|\cdot\|_{\infty}$ $\|\cdot\|_{2}$ $\|\cdot\|_{\infty}$ $\|\cdot\|_{2}$ $\|\cdot\|_{\infty}$ $\|\cdot\|_{2}$ $\|\cdot\|_{\infty}$ $\|\cdot\|_{2}$
2 0.333 / 0.366 1.000 / 1.000 0.333 0.466 0.433 0.400 0.466 0.333 0.366 0.400 0.500 0.200 0.333 0.133 0.333 0.133 0.333
@@
The MB feature extractor is based on the central moments
$\mu_{pq}\coloneqq\int_{\mathbb{R}^{2}}(x_{1}-\bar{x}_{1})^{p}(x_{2}-\bar{x}_{2})^{q}\mathop{}\!\mathrm{d}\mu(\mathbf{x}),\quad p,q\in\mathbb{N},$
where $(\bar{x}_{1},\bar{x}_{2})^{\top}\coloneqq\int_{\mathbb{R}^{2}}\mathbf{x}\mathop{}\!\mathrm{d}\mu(\mathbf{x})$ denotes the mean. As proposed in [47], we use the four features from [47, Eqs. (23), (26), (28), (31)] that are invariant under affine transformations and well-defined for $\mu\in\mathcal{P}_{\mathrm{c}}^{*}(\mathbb{R}^{2})$.
For an absolutely continuous reference $\sigma\in\mathcal{P}_{2}(\mathbb{R}^{2})$, LOT maps $\mu\in\mathcal{P}_{2}(\mathbb{R}^{2})$ to the optimal transport map $T_{\sigma}^{\mu}$ realizing the Wasserstein-2 distance
$W_{2}^{2}(\sigma,\mu)\coloneqq\min_{T_{\#}\sigma=\mu}\int_{\mathbb{R}^{2}}\|T(\mathbf{x})-\mathbf{x}\|^{2}\mathop{}\!\mathrm{d}\sigma(\mathbf{x}),$ (12)
where we optimize over all measurable $T\colon\mathbb{R}^{2}\to\mathbb{R}^{2}$. For numerical realization, instead of solving (12), we compute the optimal transport plan $\pi_{\sigma}^{\mu}$ from the Kantorovich problem
$\pi_{\sigma}^{\mu}\in\operatorname*{arg\,min}_{\pi\in\Pi(\sigma,\mu)}\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}}\|\mathbf{x}-\mathbf{y}\|^{2}\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y}),$ (13)
where $\Pi(\sigma,\mu)$ is the set of measures in $\mathcal{P}(\mathbb{R}^{2}\times\mathbb{R}^{2})$ with marginals $\sigma$ and $\mu$. Afterwards, we approximate $T_{\sigma}^{\mu}$ by the barycentric projection
$T_{\pi_{\sigma}^{\mu}}(\mathbf{x})\coloneqq\int_{\mathbb{R}^{2}}\mathbf{y}\mathop{}\!\mathrm{d}\pi^{\mu}_{\mathbf{x}}(\mathbf{y}),\quad\mathbf{x}\in\mathbb{R}^{2},$
with $\pi_{\sigma}^{\mu}(A\times B)=\int_{A}\pi^{\mu}_{\mathbf{x}}(B)\mathop{}\!\mathrm{d}\sigma(\mathbf{x})$ for any Borel sets $A,B\in\mathcal{B}(\mathbb{R}^{2})$.
For illustration, the $h_{\mathrm{}}$NR-CDTs of the entire dataset are depicted in Figure 6, where the thick lines correspond to the templates and the thin lines to the transformed images. As predicted by our theory, the lines per class are nearly perfectly aligned. This also holds for the $\mathrm{tv}$NR-CDT, which in theory requires more regularity of the measures. An inspection of Figure 7, however, reveals that the corresponding NR-CDTs of the templates are rather smooth.
Our classification results are shown in Table 2 for varying numbers of equispaced angles for the underlying Radon transform. We observe that the $h_{\mathrm{}}$NR-CDT feature representations clearly outperform the R-CDT, LOT, and Euclidean baseline and, remarkably, the classification is already perfect for a small number of chosen angles. This is also the case for MB. However, since only four features are employed, we will see that MB performs worse for more complex datasets. As the different $h_{\mathrm{}}$NR-CDT representations perform nearly on par, we henceforth consider only $\mathrm{m}$NR-CDT, $h_{\mathrm{d}}$NR-CDT and $\mathrm{tv}$NR-CDT. Moreover, in most cases, $\|\cdot\|_{2}$ outperforms $\|\cdot\|_{\infty}$ so that from now on we restrict ourselves to the Euclidean norm $\|\cdot\|_{2}$.
The computation times for this first academic experiment are reported in Table 3. While the four-dimensional MB feature extractor can be computed in less than one second, LOT is not tractable and we have to downscale the images to $64\times 64$ pixels. However, even in this case, the computation of the optimal transport plans in (13) takes around 2000 seconds. For our proposed NR-CDT approach, computing the R-CDT is the dominating component. Thereafter, NT classification based on NR-CDT features is up to 100-times faster than for R-CDT. The computational cost of the normalization procedure is negligible.
angles computation of R-CDT subsequent normalization NT classification
R-CDT $h_{\mathrm{}}$NR-CDT
2 2.0687 0.0015 0.0012 0.0019
4 4.5088 0.0014 0.0041 0.0016
8 11.9253 0.0016 0.0025 0.0012
16 22.0266 0.0021 0.0065 0.0013
32 41.9814 0.0031 0.0109 0.0021
64 84.8437 0.0048 0.0261 0.0023
128 172.2234 0.0065 0.0458 0.0021
256 322.4234 0.0133 0.1372 0.0019
Table 3: Runtime study (in sec.) for academic dataset.
To study the robustness of the NT classification we consider the polygon dataset with $10$ samples per class, which includes small non-affine perturbations. Hence, a perfect classification according to Theorems 1 and 2 cannot be expected. The accuracy results are reported in Table 4. While the Euclidean baseline and R-CDT are as bad as random guessing, $\mathrm{m}$NR-CDT and $h_{\mathrm{d}}$NR-CDT perform significantly better and, for sufficiently many Radon angles, $\mathrm{tv}$NR-CDT even reaches perfect results. This indicates its potential in real-world applications and may be attributed to the fact that $\mathrm{tv}$NR-CDT uses the whole information over the entire range of angles in $0,2\pi)$, while $\mathrm{m}$NR-CDT and $h_{\mathrm{d}}$NR-CDT only consider the largest and smallest angle information. The performance of MB deteriorates compared to the academic dataset due to the slight perturbations in the templates and lags behind $\mathrm{tv}$NR-CDT.
angles
Eucl.
MB
R-CDT $\mathrm{m}$NR-CDT $h_{\mathrm{d}}$NR-CDT $\mathrm{tv}$NR-CDT
2 0.144 0.933 0.122 0.089 0.122 0.122
@@
Towards a real-world application, we consider the historic watermark dataset. Different from the academic and the polygon dataset, the samples in each class of this dataset are visually similar but do not originate from one template measure under affine transformations. Therefore, the assumptions of Theorems 1 and 2 are violated and perfect separability is not expected. For classification, we select one sample from each class as the template measure and perform NT classification, as before. The accuracy results are reported in Table 5. We observe that all our NR-CDT variants clearly outperform the Euclidean baseline, R-CDT and MB, with $\mathrm{tv}$NR-CDT reaching an accuracy of $92\%$ when using at least 512 Radon angles.
angles
Eucl.
MB
R-CDT $\mathrm{m}$NR-CDT $h_{\mathrm{d}}$NR-CDT $\mathrm{tv}$NR-CDT
2 $0.078$ $0.406$ $0.140$ $0.516$ $0.203$ $0.200$
4 $0.219$ $\bf{0.781}$ $0.703$ $0.719$
8 $0.188$ $\bf{0.828}$ $0.782$ $0.766$
16 $0.203$ $\bf{0.859}$ $0.781$ $0.766$
32 $0.203$ $\bf{0.859}$ $0.800$ $0.822$
64 $0.203$ $0.859$ $0.813$ $\bf{0.891}$
128 $0.203$ $0.859$ $0.813$ $\bf{0.891}$
256 $0.203$ $0.859$ $0.813$ $\bf{0.891}$
512 $0.203$ $0.859$ $0.813$ $\bf{0.922}$
1024 $0.203$ $0.859$ $0.796$ $\bf{0.938}$
Table 5: NT classification accuracies for the historic watermark dataset.